OpenStudy (anonymous):

What is the area of a circle with a radius of 2 meters? Use twenty-two sevenths for pi

OpenStudy (anonymous):

@Protien

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Hero

hero (hero):

\[A_{\circ} = \pi r^2\]\[A_{\circ} = \frac{22}{7}(2)^2\]

OpenStudy (anonymous):

so what is the answer

hero (hero):

Where's your calc?

OpenStudy (anonymous):

do i need scientific or regular cal to do this

hero (hero):

Either would work for this computation.

OpenStudy (anonymous):

ok

hero (hero):

Usually, though, if you use a fraction for pi, it means to put the result in fraction form.

hero (hero):

The computation above simplifies to \[A_\circ \approx \frac{88}{7} \] But 88/7 is a improper fraction so you'll have to write it as a mixed fraction.

OpenStudy (anonymous):

so would that be 25 and one seventh?

hero (hero):

Um...no

hero (hero):

I'll help you out a bit... \[A_\circ \approx \frac{88}{7} = \frac{84 + 4}{7} = \frac{84}{7} + \frac{4}{7}\]

hero (hero):

There's a particular reason why I did that. Maybe you can figure it out.